OpenStudy (anonymous):

What is the equation of the line in slope-intercept form? the line perpendicular to y = 1/3x + 5 through (2, 1) A. y = 1/3x + 7 B. y = -1/3x + 7 C. y = 3x + 7 D. y = -3x + 7

OpenStudy (anonymous):

@alexwright 15 for me 6 for you

OpenStudy (anonymous):

B is wrong

OpenStudy (anonymous):

@geerky42

OpenStudy (anonymous):

@DJ♬Øϟϟ¥

geerky42 (geerky42):

You know what slope of that line?

geerky42 (geerky42):

only one option has correct slope

OpenStudy (anonymous):

Hm

OpenStudy (anonymous):

The slope is 1/3 lol so the answer is A

geerky42 (geerky42):

let us say slope of original line is m. So slope of line perpendicular to that original line should be \(-\dfrac{1}{m}\)

geerky42 (geerky42):

original slope is 1/3, but that's not slope of perpendicular line.

OpenStudy (anonymous):

Ohhh ok

OpenStudy (anonymous):

You my good sir, deserve a medal

geerky42 (geerky42):

ok, so m = 3. do you know what the new slope is?

geerky42 (geerky42):

m = 1/3*

OpenStudy (anonymous):

so that would be A? @geerky42

geerky42 (geerky42):

none

OpenStudy (anonymous):

That's not a choice one of these has to be right

geerky42 (geerky42):

slope of line that is perpendicular to line with slope m is \(-\dfrac{1}{m}\)

geerky42 (geerky42):

so basically you flip the fraction and make it negative.

OpenStudy (anonymous):

So it would be B

OpenStudy (anonymous):

Right? :P

OpenStudy (anonymous):

i think d is right ans. because if two lines are perpendicular then always the product of their slops =-1

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

Oh crap I typed B D:

geerky42 (geerky42):

you made it negative, but you didn't flip the fraction :P

OpenStudy (anonymous):

No B is wrong

geerky42 (geerky42):

yep. D would be the answer.

geerky42 (geerky42):

@HeyoKai

OpenStudy (anonymous):

Finally! lol thanks geerky42 you're the best ! ^_^

geerky42 (geerky42):

no problem!

OpenStudy (anonymous):

lol she said you tried

OpenStudy (anonymous):

Lol

geerky42 (geerky42):

he*

geerky42 (geerky42):

I'm male lol

OpenStudy (anonymous):

Lol xD

OpenStudy (anonymous):

lol he sorry it's the picture

OpenStudy (anonymous):